Ok so here is the question: A field is 40 feet wide and 62 feet long. How much will it cost to put a fence around this field if fencing is $1.55 per foot?

Answer by macston(5194)

L=length=62 feet; W=width=40 feet; P=perimeter
P=2(L+W)
P=2(62 feet+40 feet)
P=2(102 feet)=204 feet You would need 204 feet of fence.
(204 feet)($1.55/foot}=$316.20 ANSWER: The fencing would cost $316.20.